HIGH-RESOLUTION SPECTROSCOPY OF NITROUS ACID (HONO) AND ITS DEUTERATED SPECIES (DONO) IN THE FAR- AND MID-INFRARED REGIONS

A. DEHAYEM-KAMADJEU, J. ORPHAL, A. PERRIN, N. IBRAHIM, I. KLEINER, and J.- M. FLAUD, Laboratoire Interuniversitaire des Systèmes Atmosphériques (LISA), CNRS and Universités Paris-7 et -12, Créteil, France.

Nitrous acid (HONO) is an important source for OH radicals in tropospheric photochemistry.^{a} Currently it is measured in the atmosphere using its ultraviolet absorption bands^{b} but it has also been detected in the mid-infrared spectral region using a tuneable diode-laser.^{c} In order to complete and extend previous laboratory work on the infrared fundamental bands of trans-HONO^{d} and cis-HONO^{e} and also of its deuterated species (DONO)^{f} we have carried out a number of new spectroscopic studies, focusing on

• the first high-resolution spectra of trans- and cis-HONO and -DONO in the far-infrared region (30–120 cm^{−1}),

• the first high-resolution spectra and analysis of the ν_{1} bands of trans- and cis-DONO around 2600 cm^{−1},

• new high-resolution spectra and analysis of the Coriolis-interacting ν_{5} and ν_{6} bands of HONO around 600 cm^{−1}.^{g}

In this talk we will present the main results and some comparisons with previous studies.

aB. J. Finlayson-Pitts, J. N. Pitts, Jr., Chemistry of the Upper and Lower Atmosphere, Academic Press, 2000.

bJ. Stutz, E. S. Kim, U. Platt, P. Bruno, C. Perrino, and A. Febo, J. Geophys. Res. D 105, 14585-14592, 2000.

cC. L. Schiller, S. Locquiao, T. J. Johnson, and G. W. Harris, J. Atm. Chem. 40, 275-292, 2001.

dJ.-M. Guilmot, M. Godefroid, and M. Herman, J. Mol. Spectrosc. 160, 387-400, 1993.

eJ.-M. Guilmot, M. Godefroid, and M. Herman, J. Mol. Spectrosc. 160, 401-410, 1993.

fL. O. Halonen, C. M. Deeley, I. M. Mills and V.-M. Horneman, Can. J. Phys. 62, 1300-1305, 1984.

gC. M.Deeley, I. M. Mills, L. O. Halonen, and J. Kauppinen, Can. J. Phys. 63, 962-965, 1985.